Find the value of a and b so that (x+2) and (x-2) are the factors of p(x)=ax4+2x3-3x2+bx-4.
1
Expert's answer
2012-06-12T10:46:29-0400
(x + 2) * (x - 2) = x^2 - 4

ax^4 + 2x^3 - 3x^2 + bx - 4 | x^2 -
4
ax^4 - 4ax^2 | ax^2 + 2x + (4a-3)

2x^3 + (-3+4a)x^2
2x^3 - 8x
(4a-3)x^2
+ (b+8)x
(4a-3)x^2 +
(-16a+12)
(b+8)x - 4 + 16a - 12 =
0

{ b + 8 = 0, -4 + 16a -12 = 0 }, { a = 1, b = -8 }